OpenStudy (anonymous):

If g(x)=5x-3, evaluate g(3)

OpenStudy (anonymous):

You would replace the x in the problem with 3.

OpenStudy (anonymous):

g(3)=5(3)-3 g(3)=12

OpenStudy (anonymous):

So: 5(3)-3

OpenStudy (anonymous):

5(3) = 15 -3 = 12 The answer is 12